The world’s strongest beer, which contains 13.2% alcohol, can be found in Kulmbach, West Germany. The world’s weakest beer, which contains 0.2% alcohol, can also be found in Germany. How many liters of the world’s weakest and strongest beers are needed to make 100 liters of a mixture that is 8% alcohol?
Answer by ptaylor(2198)

Let x=amount of 13.2% beer that's needed
Then 100-x=amount of 0.2% beer that's needed
Amount of pure alcohol before they are mixed=amount of pure alcohol after they are mixed, so:
0.132x +0.002(100-x)=0.08*100
0.132x+0.2-0.002x=8 subtract 0.2 from each side and collect like terms
0.130x=7.8 divide each side by 0.130
x=60 liters amount of strong beer that's needed
100-x=100-60=40 liters----amount of weak beer needed
CK
0.132*60 +0.002*40=0.08*100
7.92+0.08=8
8=8

Hope this helps----ptaylor